A spaceship starts from Earth and accelerates with a constant acceleration of 9.8 m/s2. This is very convenient because crew and passengers on the spaceship will feel the same weight as on Earth when traveling in outer space. After how many days will an observer on Earth see the spaceship move with 15 percent of lightspeed? (Note that at this speed relativistic effects can be ignored.)

A: 1.1

B: 16

C: 32

D: 53
Answer #1

15% of the speed of light =15*3*108/100=45*106 m/s

using equation of the motion

V=U+at ...1

here U=0, V=45*106m/s, a=9.8m/s2

put these value in equation 1

45*106=0+9.8t

t=45*106/9.8

t=4591836.74 sec

t=4591836.74/24*60*60

t=53.15

t=53 so D answer